2.A rotary type compressor compresses air from 98kPaa, 27C through a compression ratio of 5 and compression efficiency of 75%, find the compressor work.

3.In an ideal Vapor Compression Refrigeration Cycle,using h2=382 kJ/kg, h3=243.15 kJ/kg, h4=243.15 kJ/kg and h1=338.14 kJ/kg, determine the refrigerating effect in kJ/kg and coefficient of performance.
Answer #1

Problem 3 is answered here.

​​​​​​Given:

  • h1= 338.14 KJ/Kg
  • h2= 382 KJ/Kg
  • h3= 243.15 KJ/Kg
  • h4= 243.15 KJ/Kg
  • Ideal vapor compression refrigeration cycle

Solution:

i) Refrigeration effect, RE:

We know that, for ideal vapor compression refrigeration cycle, Refrigeration effect, RE= (h1 - h4)

  • RE= h1 - h4
  • RE= 338.14 - 243.15
  • RE= 94.99 KJ/Kg
  • Refrigeration effect, RE= 94.99 KJ/Kg

ii) Coefficient of performance, COP:

We know that, for ideal vapor compression refrigeration cycle, Coefficient of performance, COP= (h1 - h4)/(h2 - h1)

  • COP=  \frac{h1-h4}{h2-h1}
  • COP=  \frac{338.14 - 243.15}{382-338.14}
  • COP=  \frac{94.99}{43.86}
  • COP= 2.166
  • ​​​​​​​Coefficient of performance, COP= 2.166